# EtCon: Edit-then-Consolidate for Reliable Knowledge Editing

### v1

@@ A.1 Additional Implementation Details
For our baseline experiments, we utilize the EasyEdit framework. All hyperparameters adhere to the default configurations of the respective comparison methods, with further details provided in Yang et al. (2025b); Qi et al. (2025). For our proposed EtCon method, we update only the FFN down-projection layers (mlp.down_proj) in layers 7–11 of Llama-3-8B-Instruct and layers 5–9 of Qwen2.5-7B-Instruct, following prior works Zhang et al. (2024a); Meng et al. (2022a); Geva et al. (2021); Qi et al. (2024) on knowledge editing. In the editing stage, we use AdamW with learning rate and set of TPSFT, using an edit batch size of 1 to perform 1,000 sequential single-sample edits on the same model instance. TPSFT is trained for 5 epochs with at most 6 update steps per edit using early stopping. In the consolidation stage, we optimize the inference-time policy with Group Relative Policy Optimization (GRPO). The comprehensive reward function in Equation (5) uses the following weight coefficients: for , for , for , and for . These weights were determined through extensive empirical experiments to balance factual accuracy with output quality. All specific hyperparameters are available in Table 5.
@@
| Validation Frequency | Every 5 episodes |
@@
A.2 Effect of Edited Layers on EtCon
| Layers | Reli. | Gener. | Local. | |
Llama-3-8b-Instruct |
@@ A.2 Effect of Edited Layers on EtCon
We further investigate the impact of editing different FFN layers on EtCon’s performance through an ablation study, as shown in Table 6 and Figure 6. We observe that under identical hyperparameter settings, editing early layers (Layers 7-11) outperforms deeper layers (Layers 17-21) in both locality and generalization. Analyzing the ”high reward, low performance” phenomenon in deeper layers shown in Figure 6, we find that editing deeper layers is more prone to triggering Reward Hacking. Combining existing mechanistic interpretability research—which suggests that shallow layers mainly store factual knowledge while deeper layers handle information integration and reasoning—we posit that this stems from a misalignment between retained knowledge and updated information. Editing only deeper layers may cause retained knowledge in shallow layers to conflict with the injected knowledge in deeper layers. Faced with this cognitive conflict, the LLM likely adopts a speculative strategy to maximize rewards. This reward hacking leads to internal model confusion, thereby causing a decline in performance.
A.3 Lifelong Editing
To evaluate the robustness of our framework in larger-scale lifelong editing scenarios, we extend the ZsRE sequential editing experiments to 3,000 single-sample edits on distinct ZsRE examples. At each checkpoint (600, 1200, 1800, 2400, and 3000 edits), we first run the knowledge consolidation phase and then evaluate the edited LLM before proceeding with further edits, as illustrated in Figure 7. Across all three metrics, EtCon exhibits graceful degradation: starting from the first 600 edits, its Reliability and Generalization remain high and decrease only moderately as the stream length triples, while Locality fluctuates within a narrow band without signs of collapse. In contrast, FT-M begins with much lower scores and deteriorates rapidly as more edits accumulate, with Reliability and Generalization approaching near-zero and Locality dropping sharply. These results indicate that EtCon can sustain thousands of sequential edits and still substantially outperform the baseline under lifelong editing scenarios.
A.4 Evaluation on Reasoning-Oriented Architectures
| Layers | Reli. | Gener. | Local. | |
Deepseek-R1-distill-Qwen-7B
@@ A.4 Evaluation on Reasoning-Oriented Architectures
To assess the robustness of our method on architectures with inherent reasoning capabilities, we extended our evaluation to DeepSeek-R1-Distill-Qwen-7B using 1,000 samples from the ZsRE. As reported in Table 7, our method maintains high efficacy on this reasoning-oriented architecture. As reported in Table 7, EtCon maintains high editing efficacy on this reasoning-oriented model. In particular, editing the shallow layers (Layers 5–9) yields the best trade-off, achieving 88.6% Reliability and 53.5% Generalization while preserving acceptable Locality (17.0%). This suggests that our Edit-then-Consolidate paradigm is compatible with the model’s intrinsic reasoning processes rather than disrupting them.
The green curve exhibits the fastest initial convergence (Steps 0-20). Since these layers are proximal to the output projection, the model can quickly maximize the reward by establishing a direct mapping to the target answer. However, similar to the phenomenon observed in Appendix A.2, this configuration plateaus at a lower reliability level. This suggests that modifying only the deep layers could lead the model to adopt a ”shortcut” strategy which may leave the model vulnerable to the cognitive conflict between retained shallow knowledge and updated deep injections. In contrast, the red curve displays a distinct ”warm-up” phase (Steps 0-15) followed by a sustained ascent to the highest performance plateau. This trajectory indicates that injecting knowledge into the shallow layers requires more optimization steps for the consolidate stage to propagate the changes through autoregressive generation pipeline. Crucially, this delay reflects a constructive consolidation process: the model is realigning its actual generation behavior with the updated parametric knowledge base. This results in convergence stability and reliability, demonstrating that our Edit-then-Consolidate paradigm is compatible with the intrinsic reasoning mechanisms of complex architectures.
A.5 Time Efficiency Analysis
@@ A.6 CoT Generation and Processing
A.7 Real-world evaluation Details
In this work, we follow the design Yang et al. (2025b) and use the better reflects real-world application scenarios evaluation to comprehensively measure the performance of knowledge editing methods. Specifically, our evaluation process consists of three key stages:
@@ A.7 Real-world evaluation Details
(3) LLM-as-a-Judge Evaluation: To achieve a scalable and objective evaluation, we introduce a more powerful Large Language Model (LLM) to act as a ”judge.” This judge model makes its decision by comprehensively considering the original question, the ground-truth answer (Target), and the full answer content from the edited model, ultimately providing a binary (correct/incorrect) judgment. The full judge prompts are provided in Fig 4 and Fig. 5
@@ A.8 Analysis of Reward Hacking Cases
Analysis of Reward Hacking Patterns: The two cases in Figures. 10 and 11 reveal distinct failure modes in the absence of proper reward design. In Figure 10, the model exhibits ”self-correction” behavior—correctly reasoning through the problem but then artificially inserting the target answer followed by an immediate correction. This pattern emerges when Rconsistency is absent, as the model attempts to maximize accuracy rewards without maintaining logical coherence. Figure 11 demonstrates ”answer hedging” where the model provides multiple answers to maximize the probability of including the correct one. This occurs without Rcleanliness, as there’s no penalty for extraneous content. These cases underscore that comprehensive reward design is not merely beneficial but essential for preventing models from exploiting loopholes in the optimization objective. The 11.0% and 15.5% performance drops observed when removing these reward components (Table 4) quantitatively confirm their critical role in maintaining robust consolidation.

### v2

@@ A.1 Additional Implementation Details
Baseline Configuration. For our baseline experiments, we utilize the EasyEdit framework. All hyperparameters adhere to the default configurations of the respective comparison methods, with further details provided in (Yang et al., 2025b; Qi et al., 2025).
Editing Stage. For our proposed EtCon method, we update only the FFN down-projection layers (mlp.down_proj) in layers 7–11 of Llama-3-8B-Instruct and layers 5–9 of Qwen2.5-7B-Instruct, following prior works (Zhang et al., 2024a; Meng et al., 2022a; Geva et al., 2021; Qi et al., 2024) on knowledge editing. Following the lifelong editing paradigm, we use AdamW with learning rate and set of TPSFT, sequentially performing edits (batch size 1) on the same model. TPSFT is trained for 5 epochs with at most 6 update steps per edit using early stopping.
Consolidation Stage. We optimize the inference-time policy with GRPO. The comprehensive reward function in Equation (5) uses the following weight coefficients: for , for , for , and for . These weights were determined through extensive empirical experiments to balance factual accuracy with output quality. All reward components are binary (0/1), with computed via exact string matching between the extracted answer and the ground-truth after normalization. All specific hyperparameters are available in Table 7.
@@
| Validation Frequency | Every 5 steps |
@@
A.2 Additional Results
A.2.1 Effect of Edited Layers on EtCon
| Layers | Reli. | Gener. | Local. |
@@ A.2.1 Effect of Edited Layers on EtCon
Experimental Setup. We further investigate the impact of editing different FFN layers on EtCon’s performance through an ablation study, as shown in Table 8 and Figure 7.
Main Finding. We observe that under identical hyperparameter settings, editing early layers (Layers 7-11) outperforms deeper layers (Layers 17-21) in both locality and generalization.
Reward Hacking Phenomenon. Analyzing the “high reward, low performance” phenomenon in deeper layers shown in Figure 7, we find that editing deeper layers is more prone to triggering Reward Hacking.
Mechanistic Interpretation. Existing mechanistic interpretability research suggests that shallow layers mainly store factual knowledge while deeper layers handle information integration and reasoning. Based on this, we posit that the observed phenomenon stems from a misalignment between retained knowledge and updated information.
Cognitive Conflict Analysis. Editing only deeper layers may cause retained knowledge in shallow layers to conflict with the injected knowledge in deeper layers. Faced with this cognitive conflict, the LLM likely adopts a speculative strategy to maximize rewards. This reward hacking leads to internal model confusion, thereby causing a decline in performance.
A.2.2 Portability Evaluation
Superior Cross-Architecture Portability. As shown in the Figure 8, EtCon demonstrates superior portability across architectures, achieving 26.2% on Llama-3 and 24.3% on Qwen2.5, consistently outperforming the strongest baseline FT-M (18.8% and 21.8%).
Consolidation Effectiveness. This validates that the consolidation stage effectively enables the model to utilize edited facts for reasoning.
Limitations and Future Directions. However, the absolute gains remain limited compared to single-hop reliability. We attribute this to EtCon being trained solely on single-hop QA without explicit supervision for multi-hop compositionality; thus, incorporating such supervision is a natural next step to further enhance portability.
A.3 Evaluation on Reasoning-Oriented Architectures
| Layers | Reli. | Gener. | Local. |
@@ A.3 Evaluation on Reasoning-Oriented Architectures
Experimental Setup. To assess the robustness of our method on architectures with inherent reasoning capabilities, we extended our evaluation to DeepSeek-R1-Distill-Qwen-7B using 1,000 samples from the ZsRE. Following our main experiments, we target the MLP down-projection weights (mlp.down_proj) for knowledge editing.
Shallow Layers Yield Best Trade-off. As reported in Table 9, EtCon maintains high editing efficacy on this reasoning-oriented model. In particular, editing the shallow layers (Layers 5–9) yields the best trade-off, achieving 88.6% Reliability and 53.5% Generalization while preserving acceptable Locality (17.0%). This suggests that our Edit-then-Consolidate paradigm is compatible with the model’s intrinsic reasoning processes rather than disrupting them.
Deep-Layer Shortcut Phenomenon. The green curve exhibits the fastest initial convergence (Steps 0-20). Since these layers are proximal to the output projection, the model can quickly maximize the reward by establishing a direct mapping to the target answer. However, similar to the phenomenon observed in Appendix A.2.1, this configuration plateaus at a lower reliability level. This suggests that modifying only the deep layers could lead the model to adopt a ”shortcut” strategy which may leave the model vulnerable to the cognitive conflict between retained shallow knowledge and updated deep injections.
Shallow-Layer Constructive Consolidation. In contrast, the red curve displays a distinct ”warm-up” phase (Steps 0-15) followed by a sustained ascent to the highest performance plateau. This trajectory indicates that injecting knowledge into the shallow layers requires more optimization steps for the consolidate stage to propagate the changes through autoregressive generation pipeline. Crucially, this delay reflects a constructive consolidation process: the model is realigning its actual generation behavior with the updated parametric knowledge base. This results in convergence stability and reliability, demonstrating that our Edit-then-Consolidate paradigm is compatible with the intrinsic reasoning mechanisms of complex architectures.
A.4 Time Efficiency Analysis
@@ A.5 CoT Generation and Processing
A.6 Real-world evaluation Details
In this work, we follow the design (Yang et al., 2025b) and use the better reflects real-world application scenarios evaluation to comprehensively measure the performance of knowledge editing methods. Specifically, our evaluation process consists of three key stages:
@@ A.6 Real-world evaluation Details
(3) LLM-as-a-Judge Evaluation: To achieve a scalable and objective evaluation, we introduce a more powerful Large Language Model (LLM) to act as a ”judge.” This judge model makes its decision by comprehensively considering the original question, the ground-truth answer (Target), and the full answer content from the edited model, ultimately providing a binary (correct/incorrect) judgment. The full judge prompts are provided in Figure 11 and Figure 12
A.7 Metrics for MQuAKE-CF-v2
For multi-hop task evaluation, we sample 1000 instances from MQuAKE-CF-v2 proportionally across 2-hop, 3-hop, and 4-hop questions. We report the following metrics:
Edit-wise accuracy: measures whether the model correctly answers direct queries about edited facts.
Multi-hop accuracy (2/3/4-hop): evaluates whether the model can perform multi-step reasoning that involves the edited knowledge.
@@ A.8 Analysis of Reward Hacking Cases
Analysis of Reward Hacking Patterns: The two cases in Figures. 13 and 14 reveal distinct failure modes in the absence of proper reward design. In Figure 13, the model exhibits ”self-correction” behavior—correctly reasoning through the problem but then artificially inserting the target answer followed by an immediate correction. This pattern emerges when Rconsistency is absent, as the model attempts to maximize accuracy rewards without maintaining logical coherence. Figure 14 demonstrates ”answer hedging” where the model provides multiple answers to maximize the probability of including the correct one. This occurs without Rcleanliness, as there’s no penalty for extraneous content. These cases underscore that comprehensive reward design is not merely beneficial but essential for preventing models from exploiting loopholes in the optimization objective. The 11.0% and 15.5% performance drops observed when removing these reward components (Table 5) quantitatively confirm their critical role in maintaining robust consolidation.